Provide the missing data in the following table for a distributor of martial arts products: (Round Turnover and ROI answe
Answer #1

Alpha

Sales = Turnover * Average operating assets => 5* $475,000 = $2,375,000.

Net Operaring Income = Sales * Margin =>$2,375,000*6% = $142,500.

ROI = Net Operating Income / Average operating assets  => ($142,500/$475,000) * 100 = 30.0%.

BRAVO :-

Average total assets = Sales ÷ ROI =>$332,500 ÷ 30% = $1,075,000.

Margin = Net Operating Income/ Sales => $38,700/$322,500 = 12%.

Turnover = Sales / Avg Operating assets => $322,500/$1,075,000 = 3.0 times.

CHARLIE :-

Sales = Operating Income /Margin = $52,580/11% =$478,000.

Average operating assets = Net Operating Income/ROI =>($5,580/22%) = $239,000.

Turnover = Sales / Avg Operating assets => ($478,000/$239,000) = 2.0 times.

If you have any doubts please comment on the answer.